13. 1/2 points | Previous Answers GianPSE4 22.P.055 A point charge of 3.40 nC is located at the origin and a second charge of-5.60 nC is located on the x axis at x = 1.50 m. Calculate the electric flux through a sphere centered at the origin with radius 1.00 m 384.1 Repeat the calculation for a sphere of radius 2.00 m N.m2/cHW 13+++++++++++++++ The first one is right, can you help me with the second one????? Please complete the first one to familiarize yourself with the problem. A correct answer gets a THUMBS UP!!!!!!!!!!!!!!!!!!!!!!!!!!!
Answer #1

here,

Q1 = 3.4 nC

Q1 = 3.4 * 10^-9 C

the seccond charge , Q2 = - 5.6 nC

Q2 = - 5.6 * 10^-9 C

using Gauss's law

the electric flux through a sphere of radius 2 m , phi = Qenclosed /e0

phi = ( 3.4 * 10^-9 - 5.6 * 10^-9 ) /( 8.85 * 10^-12)

phi = 248.6 N.m^2/C

the electric flux through a sphere of radius 2 m is 248.6 N.m^2 /C